Solve: 3(x – 2) – 2x = x – 6
I worked this problem out twice and the first time I got 0 and the second time I ended up with No solution. I need help because I do not know which one is correct and which one is not. Thanks
a)0
b)1
c)Identity
d)No Solution

Found 2 solutions by rmromero, tutorcecilia:
Answer by rmromero(383)


3(x – 2) – 2x = x – 6
  3x - 6 - 2x = x - 6
       x - 6  = x - 6  
            0 = 0  --> True

Absolutely the solution is not 0 and 1

There is an infinite solution for this. 

Answer by tutorcecilia(2152)

3(x – 2) – 2x = x – 6
3x-6-2x=x-6
x-6=x-6
x=x
This is an identity because x can equal any real number. Check by plugging in any real number for x and the solution will also be that number.
For example, let x=1
3(1-2)-2(1)=(1)-6
3(-1)-2=-5
-3-2=-5
-5=-5 [checks out as an identity]